# arXiv:2603.00276v1[math.OA] 27 Feb 2026

### v1

@@ Abstract
00footnotetext: 2020 Mathematics subject classification: 94A40, 81P45.
Key words: group von Neumann algebra, positive definite function, convex set, affine homeomorphism, Fourier multiplier quantum channel.
@@ 1 Introduction
Given a continuous positive definite function $\varphi\colon G\to\mathbb{C}$ with $\varphi(e)=1$, one can define, according to [KaL18, Proposition 5.4.9, p. 184], a weak* continuous unital completely positive map $M_{\varphi}\colon\mathrm{VN}(G)\to\mathrm{VN}(G)$ defined by
$M_{\varphi}(\lambda_{s})=\varphi(s)\lambda_{s},\quad s\in G.$
Such maps are the Fourier multiplier quantum channels associated with $G$. They form the core of the noncommutative harmonic analysis approach to quantum information. In [Arh25], the first author investigated Fourier multiplier channels, with a particular emphasis on finite groups, and an exhaustive description is given for concrete examples, such as the quaternion group $\mathbb{Q}_{8}$ and the dihedral group $\mathbb{D}_{4}$. As a byproduct, it is shown, remarkably, that each quantum channel acting on the matrix algebra $\mathrm{M}_{2}$ is a part of a Fourier multiplier on one of these groups.
@@
We then concentrate on compact groups. In this case, we have $\mathrm{P}_{1}(G)=\mathrm{A}(G)\cap\mathrm{P}_{1}(G)$. We equip $\mathrm{P}_{1}(G)$ with the induced topology coming from the norm topology of $\mathrm{A}(G)$, which identifies $\mathrm{P}_{1}(G)$ with the normal state space of $\mathrm{VN}(G)$. This allows us to use the general theory of facial structures of state spaces of operator algebras, in particular the correspondence between central projections and split faces developed in [AlS76], [AlS01] and [Ped18]. Our main result states that for compact groups $G$ and $H$, the von Neumann algebras $\mathrm{VN}(G)$ and $\mathrm{VN}(H)$ are *-isomorphic if and only if the convex sets $\mathrm{P}_{1}(G)$ and $\mathrm{P}_{1}(H)$ are affinely homeomorphic.
@@ 3 A description of the normal state space of a group von Neumann algebra
Recall that Eymard’s identification [Eym64] gives an isometric isomorphism $\Theta_{G}\colon\mathrm{A}(G)\to\mathrm{VN}(G)_{*}$ of Banach spaces, characterized by the property that, for every $\varphi\in\mathrm{A}(G)$, there exists a unique normal linear functional $\omega\in\mathrm{VN}(G)_{*}$ such that
@@
Recall that by definition [KaL18, Definition 2.1.5 p. 40], the Fourier-Stieltjes algebra $\mathrm{B}(G)$ of a locally compact group $G$ consists of all finite linear combinations of continuous positive definite functions. It is known that $\mathrm{A}(G)\subset\mathrm{B}(G)$, see [Pis01, p. 54]. According to [Eym64, p. 209], if $G$ is compact then we have $\mathrm{B}(G)=\mathrm{A}(G)$. In particular case, we have
@@ 4 Convex structure of $\mathrm{P}_{1}(G)$ and structure of the group von Neumann algebra
Recall that if $G$ is compact we have a *-isomorphism
@@
where $\operatorname{Irr}(G)$ is the set of all equivalence classes of irreducible unitary representations. Here $d_{\pi}$ is the degree of $\pi\in\operatorname{Irr}(G)$. We start with the following result.

### v2

# arXiv:2603.00276v2[math.OA] 03 Mar 2026

@@ Abstract
00footnotetext: 2020 Mathematics subject classification: 43A77, 43A35, 52A05, 94A40.
Key words: group von Neumann algebra, positive definite function, compact group ,convex set, affine homeomorphism, Fourier multiplier quantum channel.
@@ 1 Introduction
Given a continuous positive definite function $\varphi\colon G\to\mathbb{C}$, one can define, according to [KaL18, Proposition 5.4.9, p. 184], a weak* continuous completely positive map $M_{\varphi}\colon\mathrm{VN}(G)\to\mathrm{VN}(G)$ by
$M_{\varphi}(\lambda_{s})\overset{\mathrm{def}}{=}\varphi(s)\lambda_{s},\quad s\in G.$
This map is unital precisely when $\varphi(e)=1$. Such unital maps are called Fourier multiplier quantum channels associated with $G$. They constitute a fundamental class of quantum channels arising from noncommutative harmonic analysis. In [Arh25], the first author investigated Fourier multiplier channels, with a particular emphasis on finite groups, and an exhaustive description is given for concrete examples, such as the quaternion group $\mathbb{Q}_{8}$ and the dihedral group $\mathbb{D}_{4}$. As a byproduct, it is shown, remarkably, that each quantum channel acting on the matrix algebra $\mathrm{M}_{2}$ is a part of a Fourier multiplier on one of these groups.
@@
We then concentrate on compact groups. In this case, we have $\mathrm{P}_{1}(G)=\mathrm{A}(G)\cap\mathrm{P}_{1}(G)$. We equip $\mathrm{P}_{1}(G)$ with the induced topology coming from the norm topology of $\mathrm{A}(G)$, which identifies $\mathrm{P}_{1}(G)$ with the normal state space of $\mathrm{VN}(G)$. This allows us to use the general theory of facial structures of state spaces of operator algebras, in particular the correspondence between central projections and split faces developed in [AlS76] and [AlS01]. Our main result states that for compact groups $G$ and $H$, the von Neumann algebras $\mathrm{VN}(G)$ and $\mathrm{VN}(H)$ are *-isomorphic if and only if the convex sets $\mathrm{P}_{1}(G)$ and $\mathrm{P}_{1}(H)$ are affinely homeomorphic.
@@ 3 A description of the normal state space of a group von Neumann algebra
Let $G$ be a locally compact group equipped with a left Haar measure. According to [KaL18, Theorem 2.4.3 p. 59] and [KaL18, Lemma 2.3.6 p. 52], the Fourier algebra $\mathrm{A}(G)$ can be seen as the set of all functions $\overline{f}*\check{g}$, where $f,g\in\mathrm{L}^{2}(G)$. Here $\check{g}(s)\overset{\mathrm{def}}{=}g(s^{-1})$ for any $s\in G$. Recall that Eymard’s identification [Eym64, Théorème 3.10 p. 210] (see also [KaL18, Theorem 2.3.9 p. 54]) gives an isometric isomorphism $\Theta_{G}\colon\mathrm{A}(G)\to\mathrm{VN}(G)_{*}$ of Banach spaces, characterized by the duality bracket
(3.1) $\big\langle T,\overline{f}*\check{g}\big\rangle_{\mathrm{VN}(G),\mathrm{A}(G)}=\langle T(f),g\rangle_{\mathrm{L}^{2}(G)},\quad T\in\mathrm{VN}(G),f,g\in\mathrm{L}^{2}(G).$
Consider some function $\varphi\in\mathrm{A}(G)$. In particular, writing $\varphi=\overline{f}*\check{g}$ for some $f,g\in\mathrm{L}^{2}(G)$ we see that
[…]
$=\int_{G}\overline{(\lambda_{s^{-1}}f)(t)}g(t)\mathop{}\mathopen{}\mathrm{d}t=\int_{G}\overline{f(st)}g(t)\mathop{}\mathopen{}\mathrm{d}t=(\overline{f}*\check{g})(s)=\varphi(s),$
where we use [KaL18, p. 14] in the penultimate equality. Consequently, for every function $\varphi\in\mathrm{A}(G)$, there exists a unique normal linear functional $\omega\in\mathrm{VN}(G)_{*}$ such that
@@
Recall that by definition [KaL18, Definition 2.1.5 p. 40], the Fourier-Stieltjes algebra $\mathrm{B}(G)$ of a locally compact group $G$ consists of all finite linear combinations of continuous positive definite functions. It is known that $\mathrm{A}(G)$ is an ideal of $\mathrm{B}(G)$, see [KaL18, Proposition 2.3.3 p. 51]. According to [Eym64, p. 209], if $G$ is compact then we have $\mathrm{B}(G)=\mathrm{A}(G)$. In particular case, we have
@@ 4 Convex structure of $\mathrm{P}_{1}(G)$ and structure of the group von Neumann algebra
Recall that if $G$ is compact, there exists a *-isomorphism
@@
where $\operatorname{Irr}(G)$ is the set of all equivalence classes of irreducible unitary representations. See e.g. [Rau17, Theorem 8.3.11 p. 53]. Here $d_{\pi}$ is the degree of $\pi\in\operatorname{Irr}(G)$ and $\prod_{\pi\in\operatorname{Irr}(G)}\mathrm{M}_{d_{\pi}}$ is defined as the set $\{(x_{\pi})_{\pi\in\operatorname{Irr}(G)}:x_{\pi}\in\mathrm{M}_{d_{\pi}},\sup_{\pi\in\operatorname{Irr}(G)}\left\|x_{\pi}\right\|<\infty\}$ equipped with the canonical involution and product. We start with the following result.